OpenStudy (anonymous):

You can verify the zeros of the function y=x^2+6x-7 by using a graph and finding where the graph A. is at a minimum B. is at a maximum C. crosses the x-axis D. crosses the y-axis

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

A. x = -3 C. { x = -7, x = 1} D. y = -7

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Had to switch web browsers. LaTex no rendering properly with Safari. Set y to zero and solve for x \[x^2+6 x-7=0 \]\[(x-1) (x+7)=0 \]\[\{x=1,x=-7\} \] Take the derivative of the equation\[6+2 x=0\]and solve for x to obtain the minimum y.\[x=-3 \]The plot allows one to visually verify the algebraic results.